# Off the Planckian Locus: Using 2D Chromaticity to Improve In-Camera Color

### v2

@@ 5 Conclusion
Limitations and future work.
We note that deriving xy chromaticity still depends on the DNG optimization procedure (Supplementary Section S5). Future work could remove this dependency and explore alternative chromaticity spaces better suited for this task. Next, like other colorimetric methods, our work is limited by metamers [Logvinenko2015]—distinct spectra that produce identical camera responses. In our setting with LED illuminants, we observe increased metamers due to their narrow-band characteristics [tedla2022]. Please see Supplementary Section S1.5 for more discussion.
[…]
Supplementary Material
@@
S1.3 Polynomial/Root-Polynomial Mappings
[…]
Method Size (KB) Lightbox In-the-wild
[…]
Linear ($3\times 3$) 1.54 3.57 7.91 4.47 8.29
Poly ($3\times 9$) 3.86 3.08 9.32 4.69 10.22
Poly ($3\times 19$) 7.72 2.70 11.77 4.90 12.37
R-poly ($3\times 6$) 2.70 3.27 7.74 4.11 8.14
R-poly ($3\times 13$) 5.40 3.29 7.79 4.29 8.53
Table S2: Quantitative evaluation of polynomial and root-polynomial methods. We evaluate the performance of methods trained with 2D chromaticity input when predicting linear, polynomial (Poly), and root-polynomial (R-poly)  [polynomial, Finlayson] mappings. For both polynomial and root-polynomial mappings, we test with two different kernel sizes.
S1.4 Ablations
This section presents a series of ablation studies. We begin by comparing compare our model’s performance to baseline methods across different input types and examine performance under varying levels of noise. We also assess the impact of network architecture modifications on our proposed method. Finally, we evaluate both our model and the baselines under controlled perturbations of the white point input.
@@
Noise augmentation ablation.
Nodes and layers ablation.
White balance error ablation.
In the main paper, we assume that the ground-truth illuminant is known to isolate the performance of various CST prediction methods. In practice, however, this assumption does not hold as white-balance estimation [ccmnet, Afifi2025] modules are imperfect.
To better assess real-world performance with imperfect white points, we test our method (trained with ground truth white points) alongside several baselines on the lightbox dataset with controlled white-point offsets (Table S5). For each chart, we shift the white point by sampling a random direction and applying an offset of $0\degree$, $1\degree$, $2\degree$, or $3\degree$. Our results show that CST-MLP achieves the lowest mean angular error when the whitepoint is offset by $0\degree$$2\degree$, which falls within the typical range of white-point estimators such as CCMNet [ccmnet] and Time-Aware White Balance [Afifi2025] which achieve $1.01\degree$$2.23\degree$ mean and $0.60\degree$$1.71\degree$ median angular error. We note that when the white-point offset is large ($3\degree$), our 2D CST-MLP performs similarly to 1D CST-MLP, but still outperforms 2-CST traditional interpolation.
@@
Method Input coord. Offset: 0° Offset: 1° Offset: 2° Offset: 3°
Ang. ($\degree$) $\boldsymbol{\Delta E_{2000}}$ Ang. ($\degree$) $\boldsymbol{\Delta E_{2000}}$ Ang. ($\degree$) $\boldsymbol{\Delta E_{2000}}$ Ang. ($\degree$) $\boldsymbol{\Delta E_{2000}}$
2-CST [AdobeDNG] 1D (CCT) 4.56 9.18 5.08 9.80 6.28 11.30 7.87 13.16
3-CST [Hakki_CVPR18] 1D (CCT) 4.53 9.16 5.05 9.78 6.25 11.27 7.84 13.13
CCCNN [cccnn] 1D (CCT) 4.41 8.30 4.93 9.10 5.91 10.62 7.13 12.31
CCCNN [cccnn] 2D (xy) 4.00 8.07 4.68 8.86 5.86 10.41 7.27 12.20
EXPINV [expinv] 1D (CCT) 4.29 8.18 4.88 9.05 5.98 10.67 7.32 12.42
EXPINV [expinv] 2D (xy) 3.84 7.77 4.62 8.70 5.95 10.40 7.40 12.23
NN [nearestneighbors] 1D (CCT) 6.65 11.72 6.07 10.75 6.65 11.72 7.90 13.44
NN [nearestneighbors] 2D (xy) 4.25 8.75 5.27 9.94 7.03 12.03 9.15 14.33
CST-MLP 1D (CCT) 4.03 8.35 4.61 9.17 5.86 10.86 7.46 12.82
CST-MLP 2D (xy) 3.57 7.91 4.30 8.82 5.76 10.61 7.54 12.64
Table S5: Quantitative evaluation of methods under varying white point perturbations. We evaluate the performance of 2-CST, 3-CST, CCCNN, EXPINV, NN, and CST-MLP with 1D and 2D xy chromaticity inputs on varying degrees of white point offsets. The angular offset columns show the angular offset added to each white point during testing. White points are shifted by sampling a random direction and applying the corresponding offset. The results show that our CST MLP is the best-performing method in the $0\degree$$2\degree$ error range, and that all methods are comparable under large white-point errors ($3\degree$). Methods using 2D xy chromaticity inputs are highlighted in blue.
@@
Nodes Layers Size (KB) Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
8 1 0.41 2.70 0.78 1.68 3.47 6.31 3.24 1.03 1.98 4.15 7.56 3.83 1.32 2.52 4.91 8.67 3.80 1.30 2.47 4.93 8.64
8 2 0.69 2.65 0.80 1.69 3.39 6.20 2.93 0.87 1.91 3.84 6.76 3.69 1.28 2.41 4.65 8.30 3.70 1.26 2.42 4.70 8.38
8 3 0.97 2.66 0.81 1.71 3.45 6.14 2.96 0.89 1.89 3.81 6.80 3.63 1.25 2.42 4.59 8.28 3.70 1.26 2.45 4.65 8.46
16 1 0.79 2.73 0.86 1.73 3.45 6.33 2.93 0.87 1.84 3.80 6.64 3.65 1.19 2.41 4.64 8.30 3.68 1.24 2.43 4.70 8.46
16 2 1.85 2.65 0.83 1.72 3.41 6.08 2.88 0.87 1.85 3.80 6.59 3.59 1.19 2.41 4.62 8.04 3.64 1.20 2.45 4.59 8.23
16 3 2.91 2.64 0.85 1.68 3.34 6.00 2.88 0.85 1.83 3.75 6.59 3.60 1.24 2.39 4.58 8.20 3.61 1.24 2.38 4.53 8.21
32 1 1.54 2.66 0.82 1.70 3.41 6.14 2.89 0.87 1.85 3.75 6.67 3.60 1.22 2.38 4.60 8.18 3.60 1.20 2.43 4.56 8.23
32 2 5.66 2.63 0.83 1.70 3.38 5.98 2.94 0.84 1.87 3.77 6.79 3.56 1.17 2.35 4.64 7.99 3.60 1.19 2.40 4.54 8.28
32 3 9.79 2.65 0.84 1.72 3.40 6.00 2.85 0.84 1.80 3.70 6.56 3.73 1.25 2.39 4.56 8.21 3.68 1.24 2.43 4.53 8.16
64 1 3.04 2.66 0.85 1.72 3.42 6.09 2.89 0.86 1.86 3.80 6.78 3.58 1.20 2.45 4.63 7.98 3.60 1.21 2.48 4.58 8.12
64 2 19.29 2.62 0.82 1.64 3.34 6.11 2.89 0.83 1.80 3.78 6.74 3.55 1.26 2.39 4.54 7.88 3.64 1.24 2.44 4.49 8.05
64 3 35.54 2.65 0.82 1.68 3.38 6.06 2.89 0.85 1.82 3.75 6.74 3.63 1.20 2.39 4.57 8.13 3.52 1.20 2.37 4.44 7.87
128 1 6.04 2.64 0.84 1.68 3.37 6.09 2.88 0.83 1.83 3.76 6.74 3.57 1.19 2.42 4.59 7.90 3.64 1.23 2.43 4.49 8.21
128 2 70.54 2.64 0.84 1.63 3.34 6.14 2.88 0.84 1.81 3.73 6.69 3.56 1.24 2.44 4.55 7.83 3.56 1.23 2.45 4.51 8.00
128 3 135.04 2.67 0.81 1.69 3.41 6.10 2.89 0.85 1.82 3.79 6.67 3.56 1.25 2.35 4.55 8.03 3.55 1.25 2.42 4.46 7.85
$\boldsymbol{\Delta E_{2000}}$
8 1 0.41 6.67 3.55 5.96 8.84 11.55 7.61 4.14 6.67 9.59 13.68 8.30 5.05 7.44 10.79 14.16 7.96 4.50 7.42 10.11 13.79
8 2 0.69 6.65 3.57 5.95 8.76 11.49 7.32 4.01 6.46 9.38 12.85 8.09 4.71 7.29 10.57 13.62 7.87 4.36 7.32 10.14 13.37
8 3 0.97 6.74 3.71 6.03 8.81 11.66 7.31 4.04 6.46 9.35 12.85 8.02 4.77 7.29 10.39 13.33 7.84 4.43 7.33 10.11 13.37
16 1 0.79 6.71 3.64 5.91 8.82 11.60 7.24 3.95 6.43 9.33 12.76 8.04 4.88 7.28 10.49 13.50 7.79 4.37 7.30 10.05 13.35
16 2 1.85 6.65 3.58 6.05 8.81 11.39 7.12 3.96 6.35 9.22 12.56 7.96 4.84 7.21 10.26 13.34 7.76 4.34 7.26 10.01 12.95
16 3 2.91 6.69 3.65 5.96 8.78 11.23 7.17 4.02 6.45 9.27 12.63 7.97 4.72 7.31 10.25 13.40 7.75 4.31 7.29 10.01 12.91
32 1 1.54 6.73 3.66 6.07 8.85 11.58 7.17 3.96 6.41 9.28 12.55 7.95 4.78 7.21 10.40 13.26 7.74 4.38 7.29 10.05 13.11
32 2 5.66 6.60 3.66 5.94 8.69 11.37 7.24 4.06 6.43 9.19 12.69 7.90 4.78 7.22 10.25 13.19 7.70 4.35 7.26 9.96 13.02
32 3 9.79 6.78 3.73 6.15 8.88 11.60 7.16 3.97 6.42 9.28 12.62 8.18 4.80 7.20 10.30 13.36 7.88 4.34 7.32 9.99 12.93
64 1 3.04 6.66 3.62 5.96 8.82 11.51 7.20 3.96 6.41 9.24 12.63 7.99 4.86 7.23 10.41 13.29 7.77 4.49 7.27 9.98 13.02
64 2 19.29 6.60 3.54 5.92 8.78 11.48 7.14 3.90 6.42 9.22 12.63 7.94 4.79 7.22 10.28 13.17 7.80 4.43 7.30 9.90 12.79
64 3 35.54 6.69 3.69 6.01 8.78 11.49 7.19 4.04 6.47 9.27 12.64 8.00 4.74 7.18 10.28 13.28 7.57 4.40 7.24 9.78 12.60
128 1 6.04 6.68 3.65 6.02 8.84 11.52 7.16 3.98 6.40 9.28 12.56 7.92 4.85 7.19 10.32 13.08 7.76 4.37 7.29 9.97 12.95
128 2 70.54 6.65 3.67 5.97 8.79 11.39 7.11 3.99 6.40 9.19 12.44 7.94 4.89 7.21 10.22 13.18 7.69 4.53 7.28 9.88 12.81
128 3 135.04 6.62 3.63 5.93 8.74 11.41 7.20 3.94 6.49 9.35 12.66 7.96 4.79 7.28 10.27 13.31 7.68 4.42 7.28 9.93 12.74
S1.5 Limitations
@@
As described in the main paper, we further evaluated our method on a small in-the-wild dataset captured using mobile phones (Pixel and Samsung). We captured images with ISO level 50 on Samsung and 21 on Pixel as well as using low exposure to ensure that no color patches were clipped. Our final dataset consists of 64 images from each camera, captured under identical scenes and lighting conditions.
@@
Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
Oracle 1D 0.00 0.01 2.65 0.80 1.77 3.39 5.92 2.75 0.75 1.77 3.45 6.28
2-CST [AdobeDNG] 1D 0.07 0.01 5.04 1.60 3.95 7.11 10.94 5.09 1.75 3.95 6.94 10.95
3-CST [Hakki_CVPR18] 1D 0.11 0.01 5.03 1.61 3.92 7.07 10.94 5.11 1.75 3.98 6.92 11.03
NN [nearestneighbors] 1D 15.43 0.01 6.19 2.02 4.85 8.60 13.31 5.70 2.01 4.27 7.79 12.45
CCCNN [cccnn] 1D 68.51 13.39 4.87 1.70 3.66 6.78 10.56 5.02 1.72 3.67 6.92 10.79
EXPINV [expinv] 1D 68.53 13.39 4.91 1.71 3.67 6.83 10.49 4.77 1.46 3.44 6.74 10.67
CST-MLP 1D 1.41 0.01 4.86 1.61 3.63 6.78 10.80 4.87 1.57 3.57 6.70 10.74
NN [nearestneighbors] 2D 16.97 0.01 6.55 1.96 4.63 8.54 14.43 8.90 2.85 5.88 10.89 22.99
CCCNN [cccnn] 2D 69.01 13.48 4.63 1.57 3.44 6.42 10.08 4.51 1.62 3.42 6.26 9.63
EXPINV [expinv] 2D 69.03 13.48 4.35 1.47 3.19 6.00 9.45 4.58 1.67 3.37 6.40 9.95
CST-MLP 2D 1.54 0.01 4.47 1.47 3.43 6.19 10.08 4.46 1.48 3.23 6.01 9.94
@@
Oracle 1D 0.00 0.01 5.80 3.03 4.80 7.43 10.65 5.84 2.96 4.83 7.39 10.91
2-CST [AdobeDNG] 1D 0.07 0.01 8.79 4.69 7.47 10.97 15.64 8.82 4.75 7.49 11.01 14.87
3-CST [Hakki_CVPR18] 1D 0.11 0.01 8.79 4.68 7.44 10.97 15.50 8.84 4.79 7.52 11.01 15.04
NN [nearestneighbors] 1D 15.43 0.01 10.52 5.31 8.57 13.33 19.87 9.83 5.44 8.20 12.02 17.61
CCCNN [cccnn] 1D 68.51 13.39 8.54 4.63 7.16 10.69 15.63 8.55 4.76 7.25 10.58 15.11
EXPINV [expinv] 1D 68.53 13.39 8.50 4.56 7.18 10.49 15.60 8.35 4.55 7.01 10.60 14.71
CST-MLP 1D 1.41 0.01 8.48 4.32 7.10 10.59 15.65 8.47 4.52 6.94 10.28 15.23
NN [nearestneighbors] 2D 16.97 0.01 11.04 5.24 8.56 14.21 22.22 13.91 7.31 11.35 17.56 28.61
CCCNN [cccnn] 2D 69.01 13.48 8.19 4.32 6.80 10.41 15.19 8.03 4.23 6.56 10.49 15.37
EXPINV [expinv] 2D 69.03 13.48 7.97 4.15 6.70 10.03 15.23 8.18 4.45 7.23 10.39 14.77
CST-MLP 2D 1.54 0.01 8.29 4.29 6.85 10.61 15.30 8.16 4.45 6.74 9.90 14.58
Table S7: Qualitative evaluation of in-the-wild dataset. We compare angular error and $\Delta E_{2000}$ of traditional 2-CST/3-CST interpolation [AdobeDNG, Hakki_CVPR18], NN [nearestneighbors], CCCNN [cccnn], EXPINV [expinv], and our proposed CST-MLP. For NN, CCCNN, EXPINV, and CST-MLP, we evaluate both 1D CCT-based and 2D chromaticity-based variants (highlighted in blue). We find using 2D chromaticity coordinates leads to lower error for all methods.
@@
Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
Oracle 1D 0.00 0.01 2.38 0.61 1.33 2.73 5.30 1.59 0.55 1.10 2.06 3.40 1.37 0.48 0.96 1.79 3.00
2-CST [AdobeDNG] 1D 0.07 0.01 5.00 2.11 4.15 6.07 9.24 8.40 3.78 8.77 12.27 15.22 8.16 3.91 8.49 11.81 14.65
NN [nearestneighbors] 1D 29.53 0.01 4.46 1.61 3.20 5.65 9.24 3.07 1.13 2.11 3.81 6.47 3.27 1.16 2.21 4.16 7.23
CST-MLP 1D 1.41 0.01 3.77 1.19 2.36 4.76 8.49 2.70 0.97 1.82 3.51 5.81 2.57 0.90 1.71 3.19 5.90
NN [nearestneighbors] 2D 32.48 0.01 3.98 1.30 2.60 4.97 8.65 2.48 0.94 1.77 3.17 5.36 3.06 1.05 2.07 3.99 6.56
CST-MLP 2D 1.54 0.01 3.49 1.11 2.20 4.31 7.45 2.49 0.90 1.68 3.15 5.34 2.48 0.85 1.62 3.04 5.65
@@
Oracle 1D 0.00 0.01 4.90 2.13 3.55 5.80 9.52 3.52 1.82 2.92 4.56 6.59 3.36 1.74 2.80 4.24 6.36
2-CST [AdobeDNG] 1D 0.07 0.01 8.43 4.01 6.60 10.89 15.51 9.80 6.55 10.46 12.95 15.41 9.63 6.73 10.30 12.63 14.75
NN [nearestneighbors] 1D 29.53 0.01 7.40 3.46 5.54 9.01 14.61 5.19 2.69 4.11 6.19 9.16 5.31 2.78 4.25 6.72 9.80
CST-MLP 1D 1.41 0.01 6.30 2.76 4.42 7.40 12.78 4.55 2.37 3.57 5.39 8.08 4.42 2.29 3.48 5.49 8.35
NN [nearestneighbors] 2D 32.48 0.01 6.56 3.08 4.91 7.90 12.57 4.39 2.36 3.71 5.61 8.04 5.25 2.65 4.16 6.50 9.97
CST-MLP 2D 1.54 0.01 5.82 2.67 4.21 6.89 11.17 4.22 2.24 3.40 5.12 7.47 4.33 2.25 3.40 5.41 8.12
@@
Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
Oracle 1D 0.00 0.01 2.06 0.79 1.49 2.65 4.17 1.44 0.60 1.11 1.90 2.91 1.45 0.60 1.08 1.89 2.96
2-CST [AdobeDNG] 1D 0.07 0.01 3.88 1.76 3.49 5.38 7.12 8.14 3.64 8.51 12.11 14.49 7.87 3.66 8.26 11.56 14.01
NN [nearestneighbors] 1D 29.53 0.01 3.23 1.32 2.47 4.18 6.52 2.43 0.97 1.73 3.00 5.12 2.58 1.01 1.83 3.32 5.55
CST-MLP 1D 1.41 0.01 2.52 0.97 1.79 3.16 5.20 1.88 0.79 1.41 2.47 3.87 1.86 0.77 1.46 2.40 3.70
NN [nearestneighbors] 2D 32.48 0.01 2.83 1.10 2.10 3.67 5.69 1.93 0.82 1.46 2.44 3.91 2.25 0.91 1.65 2.83 4.65
CST-MLP 2D 1.54 0.01 2.26 0.88 1.61 2.82 4.68 1.82 0.79 1.39 2.32 3.64 1.77 0.73 1.34 2.24 3.63

### v3

@@ 5 Conclusion
Limitations. Like other colorimetric methods, our work is limited by metamers [Logvinenko2015]—distinct spectra that produce identical camera responses. In our setting with LED illuminants, we observe increased metamers due to their narrow-band characteristics [tedla2022]. Please see Supplementary Section S1.4 for more discussion.
Off the Planckian Locus: Using 2D Chromaticity to Improve In-Camera Color
@@
S1.3 Ablations
This section presents a series of ablation studies. First, we evaluate the robustness of both our model and baselines when utilizing white points predicted by an off-the-shelf illumination estimation method. Next, we compare our model’s performance to baseline methods across different input types and examine performance under varying levels of noise. Finally, we also assess the impact of network architecture modifications on our proposed method.
Utilizing white points estimated by an illumination estimation network.
We extend the white point error ablation done in Section 4 of the main paper to in-the-wild scenes by using white points derived from a cross-camera illumination estimation network, CCMNet [ccmnet]. Since this model is not fine-tuned to our camera, it represents a worst-case illumination estimation network. We directly apply it to our in-the-wild dataset to derive white points, which are then used as input to a CST method. Observing Table S2, we find that when using estimated white points, our proposed 2D parameterization outperforms 1D alternatives and traditional CCT-based interpolation. Like other CCT-based methods, our approach depends on the quality of the illumination estimate; and, aside from cases of complete white point estimation failure—where all methods exhibit large colorimetric errors (see 75th percentile in Table S2)—our method shows improved performance.
Table S2: Performance when using white points estimated by an illumination estimation method. We evaluate performance on the Pixel in-the-wild dataset using white points predicted by the off-the-shelf illumination estimation model [ccmnet]. Results are reported using angular error and $\Delta E_{2000}$. In addition to the standard mean and percentile statistics, we report a trimmed mean (Tri) where the top 10% of errors are removed to reduce the influence of scenes with large illumination estimation errors. We again observe that our proposed 2D chromaticity parametrization is advantageous when using estimated white points. Additionally, these results represent a worst-case illumination estimation scenario, as the method is cross-camera and not a camera-specific model.
Angular Error () $\boldsymbol{\Delta E_{2000}}$
Method Dim Mean Tri 25% 50% 75% Mean Tri 25% 50% 75%
2-CST [AdobeDNG] 1D 6.23 5.20 2.62 4.58 7.78 15.15 12.50 4.25 10.10 20.42
3-CST [Hakki_CVPR18] 1D 6.14 5.12 2.53 4.52 7.69 14.73 12.15 4.13 9.80 20.11
NN [nearestneighbors] 1D 7.47 6.42 3.14 5.60 9.64 19.51 15.40 5.12 12.49 25.95
CST-MLP 1D 6.09 5.10 2.50 4.30 7.84 13.40 11.02 4.16 9.19 18.10
CST-MLP(RP) 1D 6.04 4.94 2.34 4.17 7.56 11.81 9.51 3.54 7.88 15.65
NN [nearestneighbors] 2D 7.45 6.19 2.72 5.31 9.63 19.41 14.86 4.46 11.31 25.42
CST-MLP 2D 6.01 4.97 2.35 4.05 7.75 13.37 10.84 3.73 8.64 18.02
CST-MLP(RP) 2D 5.88 4.75 2.19 3.93 7.41 11.97 9.48 3.50 7.76 15.59
@@
Noise augmentation ablation.
Nodes and layers ablation.
@@
Nodes Layers Size (KB) Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90%
8 1 0.41 2.70 0.78 1.68 6.31 3.24 1.03 1.98 7.56 3.83 1.32 2.52 8.67 3.80 1.30 2.47 8.64
8 2 0.69 2.65 0.80 1.69 6.20 2.93 0.87 1.91 6.76 3.69 1.28 2.41 8.30 3.70 1.26 2.42 8.38
8 3 0.97 2.66 0.81 1.71 6.14 2.96 0.89 1.89 6.80 3.63 1.25 2.42 8.28 3.70 1.26 2.45 8.46
16 1 0.79 2.73 0.86 1.73 6.33 2.93 0.87 1.84 6.64 3.65 1.19 2.41 8.30 3.68 1.24 2.43 8.46
16 2 1.85 2.65 0.83 1.72 6.08 2.88 0.87 1.85 6.59 3.59 1.19 2.41 8.04 3.64 1.20 2.45 8.23
16 3 2.91 2.64 0.85 1.68 6.00 2.88 0.85 1.83 6.59 3.60 1.24 2.39 8.20 3.61 1.24 2.38 8.21
32 1 1.54 2.66 0.82 1.70 6.14 2.89 0.87 1.85 6.67 3.60 1.22 2.38 8.18 3.60 1.20 2.43 8.23
32 2 5.66 2.63 0.83 1.70 5.98 2.94 0.84 1.87 6.79 3.56 1.17 2.35 7.99 3.60 1.19 2.40 8.28
32 3 9.79 2.65 0.84 1.72 6.00 2.85 0.84 1.80 6.56 3.73 1.25 2.39 8.21 3.68 1.24 2.43 8.16
64 1 3.04 2.66 0.85 1.72 6.09 2.89 0.86 1.86 6.78 3.58 1.20 2.45 7.98 3.60 1.21 2.48 8.12
64 2 19.29 2.62 0.82 1.64 6.11 2.89 0.83 1.80 6.74 3.55 1.26 2.39 7.88 3.64 1.24 2.44 8.05
64 3 35.54 2.65 0.82 1.68 6.06 2.89 0.85 1.82 6.74 3.63 1.20 2.39 8.13 3.52 1.20 2.37 7.87
128 1 6.04 2.64 0.84 1.68 6.09 2.88 0.83 1.83 6.74 3.57 1.19 2.42 7.90 3.64 1.23 2.43 8.21
128 2 70.54 2.64 0.84 1.63 6.14 2.88 0.84 1.81 6.69 3.56 1.24 2.44 7.83 3.56 1.23 2.45 8.00
128 3 135.04 2.67 0.81 1.69 6.10 2.89 0.85 1.82 6.67 3.56 1.25 2.35 8.03 3.55 1.25 2.42 7.85
S1.4 Limitations
@@
As described in the main paper, we further evaluated our method on a small in-the-wild dataset captured using mobile phones (Pixel and Samsung). We captured images with ISO level 50 on Samsung and ISO 21 on Pixel as well as using low exposure to ensure that no color patches were clipped. Our final dataset comprises 150 images per camera, containing identical scenes and lighting conditions.
@@
Table S6: Quantitative evaluation of in-the-wild dataset. We compare angular error and $\Delta E_{2000}$ of traditional 2-CST/3-CST interpolation [AdobeDNG, Hakki_CVPR18], NN [nearestneighbors], CCCNN [cccnn], EXPINV [expinv], and our proposed CST-MLP. For NN, CCCNN, EXPINV, and CST-MLP, we evaluate both 1D CCT-based and 2D chromaticity-based variants (highlighted in blue). We find using 2D chromaticity coordinates leads to lower error for all methods.
@@
Mean 25% 50% 90% Mean 25% 50% 90%
Oracle 1D 0.00 0.01 1.89 0.47 1.12 4.31 2.07 0.53 1.25 4.53
2-CST [AdobeDNG] 1D 0.07 0.01 3.86 1.25 3.05 8.04 3.80 1.26 2.94 8.19
3-CST [Hakki_CVPR18] 1D 0.11 0.01 3.81 1.25 2.97 7.91 3.81 1.33 2.93 8.14
NN [nearestneighbors] 1D 15.43 0.01 5.43 1.64 4.14 11.94 5.41 1.77 3.80 12.00
CCCNN [cccnn] 1D 68.51 13.39 3.58 1.35 2.47 7.70 3.67 1.37 2.58 8.05
EXPINV [expinv] 1D 68.53 13.39 3.35 1.04 2.15 7.81 3.88 1.42 2.76 8.47
CST-MLP 1D 1.41 0.01 3.68 1.43 2.71 7.91 3.58 1.18 2.43 8.35
CST-MLP (RP) 1D 2.57 0.01 3.45 1.34 2.52 7.01 3.32 1.18 2.24 7.15
NN [nearestneighbors] 2D 16.97 0.01 5.06 1.49 3.35 11.24 5.15 1.56 3.45 11.19
CCCNN [cccnn] 2D 69.01 13.48 3.47 1.10 2.28 7.90 3.56 1.21 2.63 7.83
EXPINV [expinv] 2D 69.03 13.48 3.07 1.04 1.98 6.89 3.64 1.26 2.62 8.15
CST-MLP 2D 1.54 0.01 3.32 1.28 2.42 7.06 3.40 1.19 2.40 7.68
CST-MLP (RP) 2D 2.70 0.01 3.07 1.12 2.27 6.14 3.18 1.21 2.25 6.44
@@
Oracle 1D 0.00 0.01 4.74 2.53 3.99 8.46 4.90 2.65 4.13 8.71
2-CST [AdobeDNG] 1D 0.07 0.01 7.04 4.05 5.84 12.24 7.07 4.00 5.84 12.04
3-CST [Hakki_CVPR18] 1D 0.11 0.01 7.00 3.98 5.81 12.22 7.10 4.08 5.87 12.04
NN [nearestneighbors] 1D 15.43 0.01 9.53 5.07 7.81 18.16 9.52 5.10 7.78 17.56
CCCNN [cccnn] 1D 68.51 13.39 7.06 4.10 6.00 12.36 7.29 4.27 6.32 12.34
EXPINV [expinv] 1D 68.53 13.39 6.82 3.76 5.74 12.16 7.34 4.38 6.39 12.26
CST-MLP 1D 1.41 0.01 7.05 4.14 6.02 12.00 7.15 4.07 6.22 12.02
CST-MLP (RP) 1D 2.57 0.01 7.63 4.50 6.70 13.06 7.80 4.76 6.82 13.03
NN [nearestneighbors] 2D 16.97 0.01 9.13 4.71 7.15 17.17 9.39 4.83 7.44 17.62
CCCNN [cccnn] 2D 69.01 13.48 6.79 3.80 5.77 11.96 6.81 3.89 5.79 12.15
EXPINV [expinv] 2D 69.03 13.48 6.42 3.53 5.33 11.39 6.98 3.96 6.06 12.40
CST-MLP 2D 1.54 0.01 6.71 3.92 5.75 11.64 6.98 4.06 6.13 11.56
CST-MLP (RP) 2D 2.70 0.01 6.77 3.99 5.94 11.57 7.47 4.45 6.54 12.79
@@
Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90%
Oracle 1D 0.00 0.01 2.38 0.61 1.33 5.30 1.59 0.55 1.10 3.40 1.37 0.48 0.96 3.00
2-CST [AdobeDNG] 1D 0.07 0.01 5.00 2.11 4.15 9.24 8.40 3.78 8.77 15.22 8.16 3.91 8.49 14.65
NN [nearestneighbors] 1D 29.53 0.01 4.46 1.61 3.20 9.24 3.07 1.13 2.11 6.47 3.27 1.16 2.21 7.23
CST-MLP 1D 1.41 0.01 3.77 1.19 2.36 8.49 2.70 0.97 1.82 5.81 2.57 0.90 1.71 5.90
CST-MLP (RP) 1D 2.57 0.01 3.60 1.16 2.24 7.93 2.37 0.88 1.61 4.84 2.47 0.87 1.69 5.37
NN [nearestneighbors] 2D 32.48 0.01 3.98 1.30 2.60 8.65 2.48 0.94 1.77 5.36 3.06 1.05 2.07 6.56
CST-MLP 2D 1.54 0.01 3.49 1.11 2.20 7.45 2.49 0.90 1.68 5.34 2.48 0.85 1.62 5.65
CST-MLP (RP) 2D 2.70 0.01 3.37 1.11 2.12 7.08 2.31 0.84 1.55 4.83 2.39 0.84 1.57 5.30
@@
Oracle 1D 0.00 0.01 4.90 2.13 3.55 9.52 3.52 1.82 2.92 6.59 3.36 1.74 2.80 6.36
2-CST [AdobeDNG] 1D 0.07 0.01 8.43 4.01 6.60 15.51 9.80 6.55 10.46 15.41 9.63 6.73 10.30 14.75
NN [nearestneighbors] 1D 29.53 0.01 7.40 3.46 5.54 14.61 5.19 2.69 4.11 9.16 5.31 2.78 4.25 9.80
CST-MLP 1D 1.41 0.01 6.30 2.76 4.42 12.78 4.55 2.37 3.57 8.08 4.42 2.29 3.48 8.35
CST-MLP (RP) 1D 2.57 0.01 6.54 3.12 4.97 12.60 4.45 2.44 3.71 7.57 4.98 2.72 4.36 8.66
NN [nearestneighbors] 2D 32.48 0.01 6.56 3.08 4.91 12.57 4.39 2.36 3.71 8.04 5.25 2.65 4.16 9.97
CST-MLP 2D 1.54 0.01 5.82 2.67 4.21 11.17 4.22 2.24 3.40 7.47 4.33 2.25 3.40 8.12
CST-MLP (RP) 2D 2.70 0.01 6.12 3.06 4.75 11.31 4.50 2.43 3.77 7.78 4.89 2.65 4.25 8.44
@@
Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90%
Oracle 1D 0.00 0.01 2.06 0.79 1.49 4.17 1.44 0.60 1.11 2.91 1.45 0.60 1.08 2.96
2-CST [AdobeDNG] 1D 0.07 0.01 3.88 1.76 3.49 7.12 8.14 3.64 8.51 14.49 7.87 3.66 8.26 14.01
NN [nearestneighbors] 1D 29.53 0.01 3.23 1.32 2.47 6.52 2.43 0.97 1.73 5.12 2.58 1.01 1.83 5.55
CST-MLP 1D 1.41 0.01 2.52 0.97 1.79 5.20 1.88 0.79 1.41 3.87 1.86 0.77 1.46 3.70
CST-MLP (RP) 1D 2.57 0.01 2.35 0.94 1.68 4.68 1.74 0.75 1.34 3.39 1.78 0.74 1.38 3.62
NN [nearestneighbors] 2D 32.48 0.01 2.83 1.10 2.10 5.69 1.93 0.82 1.46 3.91 2.25 0.91 1.65 4.65
CST-MLP 2D 1.54 0.01 2.26 0.88 1.61 4.68 1.82 0.79 1.39 3.64 1.77 0.73 1.34 3.63
CST-MLP (RP) 2D 2.70 0.01 2.14 0.85 1.52 4.30 1.69 0.72 1.27 3.28 1.69 0.67 1.29 3.41